the lenght of the hypotenuse of a right triangle is 34 inches and the length of one of its legs is 16inches what is the length in inches of the other legf the right triangle
Answer by mananth(12269)

Pythagoras theorem states that
Hypotenuse ^2= leg1^2+leg2^2
Hypotenuse = 34
leg1 = 16
leg2^2= hypotenuse ^2 - leg1^2

Leg2^2= 34 ^2 - 16 ^2
Leg2^2= 1156 - 256
Leg2^2= 900
take the square root
leg2= 30 inches
m.ananth@hotmail.ca